Suppose we have the following information on the examination scores and weekly employment hours of eight students:

Exam Score      Employment Hours

      80                           10

      80                           5

      68                           15

      95                           5

      75                           21

      60                           40

      90                           0

    100                           0

Assuming Exam Score is the dependent variable (Y) and Employment Hours is the independent variable (X), calculate the simple linear regression equation by hand. Using this, test to see whether the slope on X is significant at the 0.05 level (do this by hand, as well). Make sure to show your work.

Using the model for Question 1, determine the simple linear regression equation using Excel. Using p-values, comment on the goodness of fit for this regression model.
Answer #1

Y: Exam Score = 80, 80,68,95,75,60,50, lo n=no. of Samples = 8 EY= 648 mean=ý = = 81 8 Ey2 = 53774 X: Employment = 10,8,15,51

Y = a + b Ib= -0.875 a = 9-bx = 81- (-0-875) (12) = 81+10.5 =915 Y=91.5- 0.875 x on X (6) To test to see whether the slope Si

- 0. 0.05196 VI VC) = (318-2493) = 53.on155 11265 318 - 2493 8-2 1264 TV 6) = 0.204848 ; 2 = b-b = -0.875-0 0. 204848 – – 4.2